OpenStudy (anonymous):

Please help me with this ??? 1. Write the system [ 6a +5b-5c=6 as a matrix equation [-7a+7b+4c=6 [-7a-4b-9c=-1 then identify the coefficient matrix the variable matrix and the constant matrix. A. [6 5 -5] [a] [6 ] [-7 7 4][b] = [6 ] [-7 -4 -9][c] [-1] variable matrix [6 5 -5] [-7 7 4] Constant matrix [a] [b] [c] Coefficient matrix [6 ] [6 ] [-1] B. [6 -7 -7] [a] [6] [5 7 -4] [b] = [6] [ -7 -4 -9] [c] [-1] coefficient matrix [ 6 5 -5] [-7 7 4] [-7 -4 -9] Constant matrix [a] [b] [c] variable matrix [6] [6] [-1] C. [ 6 5 -5] [a] [6] [-7 7 4] [b]= [6] [-7 -4 -9] [c] [-1] Coefficient matrix: [6 5 -5] [-7 7 4] [-7 -4 -9] Variable matrix [a] [b] [c] constant matrix: [6] [6] [-1] D. [6 5 -5] [a] [6] [-7 7 4] [b]= [6] [-7 -4 -9] [c] [-1] variable matrix: [6 5 -5] [-7 7 4] [-7 -4 -9] coefficent matrix: [a] [b] [c] constant matrix: [6] [6] [-1]

OpenStudy (anonymous):

those were the answer choices lol

OpenStudy (anonymous):

how did you get that???

OpenStudy (unklerhaukus):

its just a different example

OpenStudy (unklerhaukus):

wait a minute , *mistake*

OpenStudy (anonymous):

oh okay i was really confused

OpenStudy (unklerhaukus):

\[x + 4y = 8\\ 2x + y = 9\] \[\left[ \begin{array}{cc}1&4\\ 2&1\end{array}\right]\left[\begin{array}{c}x\\y \end{array}\right] =\left[\begin{array}{c}8\\9 \end{array}\right]\] ________________________________________ \(\left[ \begin{array}{cc}1&4\\ 2&1\end{array}\right]\) is the coefficient matrix \(\left[\begin{array}{c}x\\y \end{array}\right]\) is the variable matrix \(\left[\begin{array}{c}8\\9 \end{array}\right]\) is the constant matrix

OpenStudy (anonymous):

oh okay so basically the y and x = those numbers

OpenStudy (unklerhaukus):

yeah the 3x3 matrix of numbers are the coefficients of the variables \(a, b\) and \(c\)

OpenStudy (anonymous):

oh okay i get it now thank you